OpenStudy (anonymous):

the sum of a number and its square is 182. find the number

OpenStudy (anonymous):

x(x+1)=182=13*14 x = 13

OpenStudy (anonymous):

i would guess and check on this one. ten is too small because \[10^2+10=110\] try something bigger.

jimthompson5910 (jim_thompson5910):

Let x be this unknown number So "the sum of a number and its square is 182" means that x+x^2 = 182, let's use this to solve for x x+x^2 = 182 x^2+x = 182 x^2+x-182 = 0 Now use the quadratic formula to solve x = (-b+-sqrt(b^2-4ac))/(2a) x = (-1+-sqrt(1^2-4(1)(-182)))/(2(1)) x = (-1+-sqrt(729))/2 x = (-1+-27)/2 x = (-1+27)/2 or x=(-1-27)/2 x = 13 or x= -14 Assuming you want positive numbers only, the answer is x = 13 So the unknown number is 13 So the sum of 13 and its square 169 is 182